OpenStudy (anonymous):

factor the trinomial... 8x^2+5x-3

OpenStudy (cwrw238):

8 = 2*2*2 -3 = 3*-1 -3 = -3 * 1 we need to combine these factors to give a product = -3 and sum = middle value + 5

OpenStudy (anonymous):

yes thats what i got but then i got confused here

OpenStudy (anonymous):

\[8x(x+1)-(3(x+3)...\]

OpenStudy (cwrw238):

8 * 1 and 1 * -3 does it

OpenStudy (cwrw238):

so its (8x - 3)(x + 1 )

OpenStudy (cwrw238):

'inside' and 'outside' give + 5x